OpenStudy (nastech):

integral 1 to 2 of 1/x cubed + x squared dx

OpenStudy (mimi_x3):

Is it: \[\large \int\limits\frac{1}{x^{3}} +x^{2} dx\] ?

OpenStudy (nastech):

x cubed + x squared are all under the dividing line

OpenStudy (mimi_x3):

Is it: \[\large \int\limits\frac{1}{x^{3}+x^{2}} \] ?

OpenStudy (nastech):

yes, all of that dx

OpenStudy (mimi_x3):

Partial fractions would work..

OpenStudy (anonymous):

partial fractions are the way forward: \[\int\limits{\frac{1}{x^3 +x^2}}dx = \int\limits{\frac{1}{x^2(x+1)}dx} = \int\limits{\frac{A}{x}} + \frac{B}{x^2} + \frac{C}{x+1}dx\]

OpenStudy (anonymous):

so: \[\frac{A}{x} + \frac{B}{x^2} + \frac{C}{x+1} = \frac{1}{x^2(x+1)}\] \[Ax(x+1) + B(x+1) + Cx^2 = 1\] let x = 0 in that equation to find B let x = -1 to find C let x = 1 and sub in values of B and C to find A

OpenStudy (anonymous):

yeah i always seem to end up missing something at some point when converting to P.F

OpenStudy (experimentx):

I usually never guess PF.

OpenStudy (anonymous):

@Nastech are you ok with that?

OpenStudy (anonymous):

yo @Nastech

OpenStudy (anonymous):

i think you can also do this by writing it as x^-3 + x^2 and integrating term by term

OpenStudy (anonymous):

oh - sorry - i misread the question

OpenStudy (anonymous):

:) that would be the case if it was (1/x^3) + x^2

OpenStudy (anonymous):

yes

OpenStudy (nastech):

yes I've got the answer thanks to you all for your assistance

OpenStudy (anonymous):

awesome